# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

### v2

@@ 3. Reduced clique-cube complexes and actions on them
There exist $n\in\mathbb{N}$, for each $i\in\{1,\ldots,k\}$, a closed path $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ with $v_{i,1}=v_{i,n+1}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least once, and for any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, $l(i,j)\in\{1,\dots,n\}$ such that $v_{i,l(i,j)}=s_{i,j}$, $v_{j,l(i,j)}=t_{i,j}$ and $l(j,i)=l(i,j)$.
@@
For any $i\in\{1,\ldots,k\}$, we consider the minimum length $n_{i}$ of closed paths on $(\Gamma_{i})^{c}$ passing through every vertex at least once. Set $n=\prod_{1\leq i\leq k}n_{i}$. Then, for any $i\in\{1,\ldots,k\}$, by concatenating $\frac{n}{n_{i}}$ copies of a closed path of length $n_{i}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least once, we have a closed path $(v^{\prime}_{i,1},\ldots,v^{\prime}_{i,n},v^{\prime}_{i,n+1})$ with $v^{\prime}_{i,1}=v^{\prime}_{i,n+1}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least $\frac{n}{n_{i}}$ times (in particular, at least once).
@@
We set $(v_{1,1},\ldots,v_{1,n},v_{1,n+1}):=(v^{\prime}_{1,1},\ldots,v^{\prime}_{1,n},v^{\prime}_{1,n+1})$. For any $j\in\{2,\ldots,k\}$, we define $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$ inductively as follows. Take $j\in\{2,\ldots,k\}$. Assume that $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ is defined for $i\in\{1,\ldots,j-1\}$ with $(V_{i},V_{j})\in E(T)$. Then, we set $l(i,j)$ as the minimum $l$ such that $v_{i,l}=s_{i,j}$. By a cyclic permutation of $v^{\prime}_{j,1},\ldots,v^{\prime}_{j,n}$, we have $v_{j,1},\ldots,v_{j,n}$ such that $v_{j,l(i,j)}=t_{i,j}$. By setting $v_{j,n+1}:=v_{j,1}$, we have $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$. For any $i,j$ with $1\leq i<j\leq k$ and $(V_{i},V_{j})\in E(T)$, we set $l(j,i):=l(i,j)$. ∎
We take $n\in\mathbb{N}$, $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ for each $i\in\{1,\ldots,k\}$, and $l(i,j)\in\{1,\dots,n\}$ for any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$ as in Lemma 4.1. For $1\leq l\leq n$, set
$\lambda_{l}:=v_{1,l}v_{2,l}\cdots v_{k,l}.$
For any $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, we define $\lambda_{l}(i,j)$ as $\lambda_{l}(i,j):=\lambda_{l}$ if $l\neq l(i,j)$. In addition, we define $\lambda_{l(i,j)}(i,j)$ as
$\lambda_{l(i,j)}(i,j):=\begin{array}[]{lll}\tau_{i,j}v_{1,l(i,j)}v_{2,l(i,j)}\cdots v_{i-1,l(i,j)}v_{i+1,l(i,j)}&&\\ \ \ \cdots v_{j-1,l(i,j)}v_{j+1,l(i,j)}\cdots v_{k,l(i,j)}\end{array}\text{if }i<j,\text{ and}$
$\lambda_{l(i,j)}(i,j):=\begin{array}[]{lll}\tau_{i,j}v_{1,l(i,j)}v_{2,l(i,j)}\cdots v_{j-1,l(i,j)}v_{j+1,l(i,j)}&&\\ \ \ \cdots v_{i-1,l(i,j)}v_{i+1,l(i,j)}\cdots v_{k,l(i,j)}\end{array}\text{if }i>j,$
where we note that $l(j,i)=l(i,j)$. Moreover, we define $\gamma_{\natural}(i,j)$ as
$\gamma_{\natural}(i,j):$ $=\lambda_{1}(i,j)\lambda_{2}(i,j)\cdots\lambda_{n}(i,j)$
$=\lambda_{1}\lambda_{2}\cdots\lambda_{l(i,j)-1}\lambda_{l(i,j)}(i,j)\lambda_{l(i,j)+1}\cdots\lambda_{n}.$
Take a closed path
(4.1) $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$
with $i_{1}=i_{r+1}=1$ on the spanning tree $T$ of $Q(\Gamma,\Gamma_{0})$ passing through every vertex at least once. We define $\gamma_{\natural}$ as
(4.2) $\gamma_{\natural}:=\gamma_{\natural}(i_{1},i_{2})\gamma_{\natural}(i_{2},i_{3})\cdots\gamma_{\natural}(i_{r},i_{r+1}).$
For $h\geq 1$, we set
@@
Since $\Gamma^{t}$ is connected, we have $\mathfrak{d}\in\mathbb{N}$ such that $W(\mathfrak{d})\neq\emptyset$ and $V_{0}=W(1)\sqcup\cdots\sqcup W(\mathfrak{d})$. We define a map $V_{0}\to\{1,\ldots,k\}\times\{1,\ldots,n\};w\mapsto(i(w),l(w))$ as $(i(w),l(w))$ is the minimum of $(i,l)$ indexing $v_{i,l}\in V_{*}$ such that $d_{\Gamma^{t}}(w,v_{i,l})=\min_{v\in V_{\ast}}d_{\Gamma^{t}}(w,v)$ with respect to the dictionary order on $\{1,\ldots,k\}\times\{1,\ldots,n\}$. In particular for $w\in W(1)$, $(i(w),l(w))$ is the minimum of $(i,l)$ with $\mu(w,v_{i,l})>2$ with respect to the dictionary order on $\{1,\ldots,k\}\times\{1,\ldots,n\}$. We set
(4.4) $W(0):=\{v_{i(w),l(w)}\in V_{\ast}\mid w\in W(1)\},$
which equals $\{v_{i(w),l(w)}\in V_{\ast}\mid w\in V_{0}\}$. Note that for $w\in W(0)$, $i$ and $l$ with $w=v_{i,l}$, $i$ is unique, but $l$ is not unique. Therefore we define for $w\in W(0)$, $(i(w),l(w))$ is the minimum of $(i,l)$ with $w=v_{i,l}$ with respect to the dictionary order on $\{1,\ldots,k\}\times\{1,\ldots,n\}$.
We define a total order $<$ on $W(0)\sqcup V_{0}=W(0)\sqcup\cdots\sqcup W(\mathfrak{d})$ as follows. First we take a total order $<$ on $W(0)$ such that for any $w,w^{\prime}\in W(0)$, $(i(w),l(w))<(i(w^{\prime}),l(w^{\prime}))$ if and only if $w<w^{\prime}$. Next suppose that for $h\geq 1$, $W(0)\sqcup\cdots\sqcup W(h-1)$ is equipped with a total order $<$. Then we define a map
(4.5) $W(h)\to W(h-1);w\mapsto w(h-1)$
as $w(h-1)$ is the minimum of $w^{\prime}\in W(h-1)$ with $\mu(w,w^{\prime})>2$ with respect to the given total order $<$. We take a total order $<$ on $W(h)$ such that for any $w,w^{\prime}\in W(h)$, $w(h-1)\leq w^{\prime}(h-1)$ if $w<w^{\prime}$. Moreover for any $w\in W(0)\sqcup\cdots\sqcup W(h-1)$ and $w^{\prime}\in W(h)$, we set $w<w^{\prime}$. Then $<$ on $W(0)\sqcup\cdots\sqcup W(h-1)$ is extended on $W(0)\sqcup\cdots\sqcup W(h)$. Finally we have the order $<$ on $W(0)\sqcup V_{0}=W(0)\sqcup\cdots\sqcup W(\mathfrak{d})$. For any $0\leq h\leq\mathfrak{d}$ and $w\in W(h)$, we take the path $(w(h),w(h-1),\ldots,w(0))$ such that $w(h)=w$, $w(\mathfrak{j}-1)=w(\mathfrak{j})(\mathfrak{j}-1)$ for any $2\leq\mathfrak{j}\leq h$, and $w(0)=v_{i(w),l(w)}$, where $(i(w),l(w))=(i(w(1)),l(w(1)))$. See Figure 4.
Figure 4. $W(0),\ldots,W(\mathfrak{d})$ in the case of $\mathfrak{d}=3$.
Now we name all elements of $W(0)\sqcup\cdots\sqcup W(\mathfrak{d}-1)=W(0)\sqcup V_{0}\setminus W(\mathfrak{d})$ as
(4.6) $w_{1},\ldots,w_{\mathfrak{m}}$
such that $w_{1}<\cdots<w_{\mathfrak{m}}$ where $\mathfrak{m}:=\#(W(0)\sqcup\cdots\sqcup W(\mathfrak{d}-1))$. For any $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$, we define $\lambda_{l}(\mathfrak{j})$ as $\lambda_{l}(\mathfrak{j}):=\lambda_{l}$ if $l\neq l(w_{\mathfrak{j}})$. In addition, we take $h\in\{0,1,\ldots,\mathfrak{d}-1\}$ with $w_{\mathfrak{j}}\in W(h)$ and define $\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})$ as
$\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j}):=v_{1,l(w_{\mathfrak{j}})}\cdots v_{i(w_{\mathfrak{j}})-1,l(w_{\mathfrak{j}})}v_{i(w_{\mathfrak{j}})+1,l(w_{\mathfrak{j}})}\cdots v_{k,l(w_{\mathfrak{j}})}$
$w_{\mathfrak{j}}(h)w_{\mathfrak{j}}(h-1)\cdots w_{\mathfrak{j}}(1)w_{\mathfrak{j}}(0),$
where we note $w_{\mathfrak{j}}(h)=w_{\mathfrak{j}}$ and $w_{\mathfrak{j}}(0)=v_{i(w_{\mathfrak{j}}),l(w_{\mathfrak{j}})}$. We set
(4.7) $\gamma_{\flat}(\mathfrak{j}):=\lambda_{1}(\mathfrak{j})\cdots\lambda_{n}(\mathfrak{j})=\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}})-1}\lambda_{l(w_{\mathfrak{j}})}(\mathfrak{j})\lambda_{l(w_{\mathfrak{j}})+1}\cdots\lambda_{n}.$
We define $\gamma_{\flat}$ as
(4.8) $\gamma_{\flat}:=\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{m}).$
By using (4.8) and (4.2), we define
(4.9) $\gamma:=\gamma_{\flat}\gamma_{\natural}=\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{m})\gamma_{\natural}(i_{1},i_{2})\gamma_{\natural}(i_{2},i_{3})\cdots\gamma_{\natural}(i_{r},i_{r+1}).$
@@
$\gamma(0):=1,\gamma(1):=\lambda_{1}(1),\gamma(2):=\lambda_{1}(1)\lambda_{2}(1),\ldots,\gamma(n):=\gamma_{\flat}(1),$
for $a\in\{2,\ldots,\mathfrak{m}\}$ and $l\in\{1,\ldots,n\}$,
$\gamma((a-1)n+l):=\gamma_{\flat}(1)\cdots\gamma_{\flat}(a-1)(\lambda_{1}(a)\cdots\lambda_{l}(a)).$
In particular, we have $\gamma(\mathfrak{m}n)=\gamma_{\flat}$. Moreover we set
$\gamma(\mathfrak{m}n+1):=\gamma_{\flat}\lambda_{1}(i_{1},i_{2}),\gamma(\mathfrak{m}n+2):=\gamma_{\flat}\lambda_{1}(i_{1},i_{2})\lambda_{2}(i_{1},i_{2}),\ldots,\gamma(\mathfrak{m}(n+1)):=\gamma_{\flat}\gamma_{\natural}(i_{1},i_{2}),$
for $a\in\{2,\ldots,r\}$ and $l\in\{1,\ldots,n\}$,
$\gamma((\mathfrak{m}+a-1)n+l):=\gamma_{\flat}\gamma_{\natural}(i_{1},i_{2})\cdots\gamma_{\natural}(i_{a-1},i_{a})(\lambda_{1}(i_{a},i_{a+1})\cdots\lambda_{l}(i_{a},i_{a+1})).$
In particular, we have $\gamma((\mathfrak{m}+r)n)=\gamma$.
We confirm the following for convenience:
@@
    $k$ is the number of indecomposable factors of a unique join decomposition of $\Gamma_{\ast}$;
@@
Then, $U_{l}$ spans a clique containing $\Gamma_{0}$ in $\Gamma$, that is, $U_{l}\in\Delta_{\Gamma,\Gamma_{0}}$. Hence, we have a $k$-dimensional cube $[A_{V_{0}},A_{U_{l}}]$ in $C_{\Gamma,\Gamma_{0}}$. The hyperplanes dual to edges of $[A_{V_{0}},A_{U_{l}}]$ are of $v_{i,l}$-type ($i\in\{1,\ldots,k\}$). We denote such hyperplanes as $H_{i,l}$ ($i\in\{1,\ldots,k\}$).

### v3

@@ 3. Reduced clique-cube complexes and actions on them
There exist $n\in\mathbb{N}$, a matrix $M_{\natural}=(v_{i,l})\in\mathcal{M}(k,n;V)$, and a map
(4.7) $l_{\flat}:\left\{(i,j)\middle|\begin{array}[]{l}1\leq i,j\leq k,\\ (V_{i},V_{j})\in E(T)\end{array}\right\}\to\{1,\ldots,n\}$
which satisfy the following:
  • (i)
    For each $i\in\{1,\ldots,k\}$, $(v_{i,1},\ldots,v_{i,n},v_{i,1})$ is a closed path on $(\Gamma_{i})^{c}$ that passes through every vertex at least once;
  • (ii)
    For each $i,j\in\{1,\ldots,k\}$ with $(V_{i},V_{j})\in E(T)$, $v_{i,l_{\flat}(i,j)}=s_{i,j}$, $v_{j,l_{\flat}(i,j)}=t_{i,j}$;
  • (iii)
    $l_{\flat}({j,i})=l_{\flat}(i,j)$.
@@
For any $i\in\{1,\ldots,k\}$, we consider the length $n_{i}$ of a closed path on $(\Gamma_{i})^{c}$ passing through every vertex at least once. Set $n=\prod_{1\leq i\leq k}n_{i}$. Then, for any $i\in\{1,\ldots,k\}$, by concatenating $\frac{n}{n_{i}}$ copies of a closed path of length $n_{i}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least once, we have a closed path $(v^{\prime}_{i,1},\ldots,v^{\prime}_{i,n},v^{\prime}_{i,n+1})$ with $v^{\prime}_{i,1}=v^{\prime}_{i,n+1}$ on $(\Gamma_{i})^{c}$ passing through every vertex at least $\frac{n}{n_{i}}$ times (in particular, at least once).
@@
We set $(v_{1,1},\ldots,v_{1,n},v_{1,n+1}):=(v^{\prime}_{1,1},\ldots,v^{\prime}_{1,n},v^{\prime}_{1,n+1})$. For any $j\in\{2,\ldots,k\}$, we define $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$ inductively as follows. Take $j\in\{2,\ldots,k\}$. Suppose that $(v_{i,1},\ldots,v_{i,n},v_{i,n+1})$ is defined for $i\in\{1,\ldots,j-1\}$ with $(V_{i},V_{j})\in E(T)$. Note that such $i$ is unique for $j$. Then, we define $l_{\flat}(i,j)$ as the minimum $l$ such that $v_{i,l}=s_{i,j}$. By a cyclic permutation of $v^{\prime}_{j,1},\ldots,v^{\prime}_{j,n}$, we have $v_{j,1},\ldots,v_{j,n}$ such that $v_{j,l_{\flat}(i,j)}=t_{i,j}$. By setting $v_{j,n+1}:=v_{j,1}$, we have $(v_{j,1},\ldots,v_{j,n},v_{j,n+1})$. Finally, for any $i,j$ with $1\leq i<j\leq k$ and $(V_{i},V_{j})\in E(T)$, we set $l_{\flat}(j,i):=l_{\flat}(i,j)$. ∎
We take $n\in\mathbb{N}$, $M_{\natural}=(v_{i,l})\in\mathcal{M}(k,n;V)$ and a map $l_{\flat}$ as in Lemma 4.2. In the case where $k=1$, $E(T)=\emptyset$, and thus $l_{\flat}$ and the conditions (ii), (iii) are meaningless. Nonetheless, the condition (i) remains meaningful.
Now we define $\gamma_{\natural}\in A_{V_{\ast}}$ as
(4.8) $\gamma_{\natural}:=prod(M_{\natural})=(v_{1,1}v_{2,1}\cdots v_{k,1})\cdots(v_{1,n}v_{2,n}\cdots v_{k,n}).$
When $\Gamma_{0}=\emptyset$ and $k=1$, this $\gamma_{\natural}$ is a WPD contracting element on $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$ [14]. We do not directly need this fact. Even in the general setting, we can confirm that $\gamma_{\natural}$ fits the condition (i) of Theorem 2.9. Unfortunately, it is unclear whether $\gamma_{\natural}$ fits the conditions (ii) and (iii). To more easily approach these conditions, we prepare two modifications $\gamma_{\sharp}$ and $\gamma_{\flat}$ based on powers of $\gamma_{\natural}$.
We prepare for the definition of $\gamma_{\flat}$. Let $e$ be the identity in $A_{\Gamma}$. When $k=1$ (for example, $\Gamma$ in Figure 4), we define $\gamma_{\flat}=e$. Suppose that $k\geq 2$. Recall that we have $n\in\mathbb{N}$, $M_{\natural}=(v_{i,l})\in\mathcal{M}(k,n;V)$, a map $l_{\flat}$ as in Lemma 4.2, and a closed path $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ with $i_{1}=i_{r+1}=1$, defined in (4.6). Note that $r\geq 2$, since $k\geq 2$. We define supplementary matrices $M^{\prime}_{\flat}(a),M^{\prime\prime}_{\flat}(a)\in\mathcal{M}(k,n;A_{\Gamma})$ $(a\in\{1,\ldots,r\})$ by
(4.9) $\left(M^{\prime}_{\flat}(a)\right)[i,l]=\begin{cases}v_{i,l}^{-1}\quad&(i\in\{i_{a},i_{a+1}\},l=l_{\flat}(i_{a},i_{a+1}))\\ e\quad&(\text{otherwise}),\end{cases}$
(4.10) $\left(M^{\prime\prime}_{\flat}(a)\right)[i,l]=\begin{cases}\tau_{i_{a},i_{a+1}}\quad&(i=1,l=l_{\flat}(i_{a},i_{a+1}))\\ e\quad&(\text{otherwise}).\end{cases}$
Using $M^{\prime}_{\flat}(a)$ and $M^{\prime\prime}_{\flat}(a)$, we define $M_{\flat}(a)\in\mathcal{M}(k,n;A_{\Gamma})$ by
(4.11) $M_{\flat}(a):=M^{\prime\prime}_{\flat}(a)\odot M^{\prime}_{\flat}(a)\odot M_{\natural},$
[…]
(4.12) $M_{\flat}=cbind\left(M_{\flat}(1),\ldots,M_{\flat}(r)\right).$
For the definition of $cbind$, see (2.2).
We now define $\gamma_{\flat}\in A_{\Gamma}$ as
(4.13) $\gamma_{\flat}:=prod(M_{\flat})=prod(M_{\flat}(1))\cdots prod(M_{\flat}(r)).$
For the definition of the map $prod$, see (2.5) and (2.6). $\gamma_{\flat}$ is a modification of $\gamma_{\natural}^{r}$. When $k\geq 2$ and $\Gamma_{0}=\emptyset$, we can take $\gamma_{\flat}$ as a WPD element [29]. This fact is not used directly. To deal with the case where $\Gamma_{0}\neq\emptyset$, we need the other modification, $\gamma_{\sharp}$.
To explain the construction of $\gamma_{\sharp}$, we introduce the “height” $h$ of elements of $V_{0}$ in the case $V_{0}\neq\emptyset$. Otherwise, we define $\gamma_{\sharp}=e$. For $h\geq 1$, we set
@@
Since $\Gamma^{t}$ is connected, we have $\mathfrak{h}\in\mathbb{N}$ such that $W(\mathfrak{h})\neq\emptyset$ and $V_{0}=W(1)\sqcup\cdots\sqcup W(\mathfrak{h})$. Suppose that $\mathfrak{h}\geq 2$. Otherwise, we define $\gamma_{\sharp}=e$. We define two maps
(4.15) $i_{\sharp}:V_{0}\to\{1,\ldots,k\},$
(4.16) $l_{\sharp}:V_{0}\to\{1,\ldots,n\}$
such that $(i_{\sharp}(w),l_{\sharp}(w))$ is the minimum of $(i,l)$ indexing $v_{i,l}\in V_{*}$ with
$d_{\Gamma^{t}}(w,v_{i,l})=\min_{v\in V_{\ast}}d_{\Gamma^{t}}(w,v)$
with respect to the order $<$ on $\{1,\ldots,k\}\times\{1,\ldots,n\}$ defined by
$(i,l)<(i^{\prime},l^{\prime})\Leftrightarrow(l<l^{\prime})\text{ or }(l=l^{\prime}\text{ and }i<i^{\prime}).$
Note that this is a slightly unusual dictionary order, which considers the second factor first. In particular, for $w\in W(1)$, $(i_{\sharp}(w),l_{\sharp}(w))$ is the minimum of $(i,l)$ with $\mu(w,v_{i,l})>2$ with respect to the above order.
@@
(4.17) $W(0):=\{v_{i_{\sharp}(w),l_{\sharp}(w)}\in V_{\ast}\mid w\in W(1)\},$
which equals $\{v_{i_{\sharp}(w),l_{\sharp}(w)}\in V_{\ast}\mid w\in V_{0}\}$. Obviously, we have $W(0)\cap V_{0}=\emptyset$. We extend two maps $i_{\sharp}$ and $l_{\sharp}$ on $W(0)\sqcup V_{0}$ by defining $(i_{\sharp}(w),l_{\sharp}(w))$ for $w\in W(0)$ as the minimum of $(i,l)$ with $w=v_{i,l}$ with respect to the order $<$ on $\{1,\ldots,k\}\times\{1,\ldots,n\}$. Note that for $w\in W(0)$, $i$ and $l$ with $w=v_{i,l}$, $i$ is unique, but $l$ is not necessarily unique.
Next, by using the order $<$ on $\{1,\ldots,k\}\times\{1,\ldots,n\}$ and the two maps $i_{\sharp}$, $l_{\sharp}$ on $W(0)\sqcup V_{0}=W(0)\sqcup W(1)\sqcup\cdots\sqcup W(\mathfrak{h})$, we define a total order $\prec$ on $W(0)\sqcup V_{0}$ as follows. First, we take a total order $\prec$ on $W(0)$ such that for any $w,w^{\prime}\in W(0)$, $w\prec w^{\prime}$ if and only if $(i_{\sharp}(w),l_{\sharp}(w))<(i_{\sharp}(w^{\prime}),l_{\sharp}(w^{\prime}))$. Next, suppose that for $1\leq h\leq\mathfrak{h}$, $W(0)\sqcup\cdots\sqcup W(h-1)$ is equipped with a total order $\prec$. Then we define a map
$\pi_{h}:W(h)\to W(h-1)$
as $\pi_{h}(w)$ $(w\in W(h))$ is the minimum of $w^{\prime}\in W(h-1)$ with $\mu(w,w^{\prime})>2$ with respect to the given total order $\prec$. We take a total order $\prec$ on $W(h)$ such that for any different $w,w^{\prime}\in W(h)$, $w\prec w^{\prime}$ only if $\pi(w)\preceq\pi(w^{\prime})$. Moreover, for any $w\in W(0)\sqcup\cdots\sqcup W(h-1)$ and $w^{\prime}\in W(h)$, we set $w\prec w^{\prime}$. Then $\prec$ on $W(0)\sqcup\cdots\sqcup W(h-1)$ is extended on $W(0)\sqcup\cdots\sqcup W(h)$. Finally, we have the order $\prec$ on $W(0)\sqcup V_{0}=W(0)\sqcup\cdots\sqcup W(\mathfrak{h})$. We define a map
(4.18) $\pi:W(1)\sqcup\cdots\sqcup W(\mathfrak{h})\to W(0)\sqcup\cdots\sqcup W(\mathfrak{h}-1),$
by $\pi(w)=\pi_{h}(w)\in W(h-1)$ for $w\in W(h)$. For any $w\in W(0)\sqcup V_{0}$, let $h\in\{0,\ldots,\mathfrak{h}\}$ be the height of $w$; that is, $w\in W(h)$. We take a path from $w$ to $v_{i_{\sharp}(w),l_{\sharp}(w)}$ on $\Gamma^{t}$;
(4.19) $(w(h),w(h-1),\ldots,w(0)):=(w,\pi(w),\pi^{2}(w),\ldots,\pi^{h}(w)).$
We note that $(i_{\sharp}(w),l_{\sharp}(w))=(i_{\sharp}(w(h^{\prime})),l_{\sharp}(w(h^{\prime})))$ for any $0\leq h^{\prime}\leq h$. In particular, $w(0)=\pi^{h}(w)=v_{i_{\sharp}(w),l_{\sharp}(w)}$. See Figure 7.
Figure 7. In the figure, $w$ is a vertex in $W(1)$, and thin solid lines indicate edges labeled by $2$. A thick solid line, which connects $w$ and $v_{i_{\sharp}(w),l_{\sharp}(w)}$, indicates an edge labeled by an integer greater than $2$. For all edges which are presented as dotted lines, we do not care about their label.
Now we name all elements of $W(1)\sqcup\cdots\sqcup W(\mathfrak{h}-1)=V_{0}\setminus W(\mathfrak{h})$ as
(4.20) $w_{1},w_{2},\ldots,w_{\mathfrak{m}}$
such that $w_{1}\prec w_{2}\prec\cdots\prec w_{\mathfrak{m}}$ where
(4.21) $\mathfrak{m}:=\#(W(1)\sqcup\cdots\sqcup W(\mathfrak{h}-1)).$
Moreover, we set the heights $h_{1},h_{2},\ldots,h_{\mathfrak{m}}$ for $w_{1},w_{2},\ldots,w_{\mathfrak{m}}$, respectively. We have $11=h_{1}\leq h_{2}\leq\cdots\leq h_{\mathfrak{m}}=\mathfrak{h}-$ 1 . The following is clear by the construction of the order $\prec$ and the map $\pi$.
Lemma 4.3.
Let $h\in\{1,\ldots,\mathfrak{h}\}$ and $w\in W(h)$.
@@
    Let $h=1$. Let $w^{\prime}\in V_{0}$ with $w^{\prime}\succeq w$. For any $(i,l)\in\{1,\ldots,k\}\times\{1,\ldots,n\}$ with $(i,l)<(i_{\sharp}(w),l_{\sharp}(w))$, we have $\mu(v_{i,l},w^{\prime})=2$.
  2. (2)
    Let $h\geq 2$. Let $w^{\prime}\in V_{0}$ with $w^{\prime}\succeq w$.
    1. (i)
      For any $v\in V_{\ast}$, we have $\mu(v,w^{\prime})=2$.
    2. (ii)
      For any $w^{\prime\prime}\in V_{0}$ with $w^{\prime\prime}\prec\pi(w)$, we have $\mu(w^{\prime\prime},w^{\prime})=2$.
See Figure 7.
We now explain the construction of $\gamma_{\sharp}$. If $V_{0}=\emptyset$ or $\mathfrak{h}=1$, then we have already defined $\gamma_{\sharp}=e$. Suppose that $V_{0}\neq\emptyset$ and that $\mathfrak{h}\geq 2$. We define supplementary matrices $M^{\prime}_{\sharp}(\mathfrak{j}),M^{\prime\prime}_{\sharp}(\mathfrak{j})\in\mathcal{M}(k,n;A_{\Gamma})$ $(\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\})$ by
(4.22) $\left(M^{\prime}_{\sharp}(\mathfrak{j})\right)[i,l]=\begin{cases}v_{i,l}^{-1}&\quad(i=i_{\sharp}(w_{\mathfrak{j}}),l=l_{\sharp}(w_{\mathfrak{j}}))\\ e&\quad(\text{otherwise})\end{cases}$
and
(4.23) $\left(M^{\prime\prime}_{\sharp}(\mathfrak{j})\right)[i,l]=\begin{cases}w_{\mathfrak{j}}\left(h_{\mathfrak{j}}\right)w_{\mathfrak{j}}\left(h_{\mathfrak{j}}-1\right)\cdots w_{\mathfrak{j}}(0)&\quad(i=k,l=l_{\sharp}(w_{\mathfrak{j}}))\\ e&\quad(\text{otherwise}),\end{cases}$
where $h_{\mathfrak{j}}$ is the height of $w_{\mathfrak{j}}$. Using $M^{\prime}_{\sharp}(\mathfrak{j})$ and $M^{\prime\prime}_{\sharp}(\mathfrak{j})$, we define $M_{\sharp}(\mathfrak{j})\in\mathcal{M}(k,n;A_{\Gamma})$ by
(4.24) $M_{\sharp}(\mathfrak{j}):=M_{\natural}\odot M^{\prime}_{\sharp}(\mathfrak{j})\odot M^{\prime\prime}_{\sharp}(\mathfrak{j}),$
[…]
(4.25) $M_{\sharp}=cbind\left(M_{\sharp}(1),\ldots,M_{\sharp}(\mathfrak{m})\right).$
For the definition of $cbind$, see (2.2). We now define $\gamma_{\sharp}\in A_{\Gamma}$ as
(4.26) $\gamma_{\sharp}:=prod(M_{\sharp})=prod(M_{\sharp}(1))\cdots prod(M_{\sharp}(\mathfrak{m})).$
$\gamma_{\sharp}$ is a modification of $\gamma_{\natural}^{\mathfrak{m}}$.
We define $M_{\natural\sharp\flat}\in\mathcal{M}(k,(1+\mathfrak{m}+r)n;A_{\Gamma})$ by
(4.27) $M_{\natural\sharp\flat}:=cbind(M_{\natural},M_{\sharp},M_{\flat}).$
Now, we define $\gamma_{\natural\sharp\flat}\in A_{\Gamma}$ by
(4.28) $\gamma_{\natural\sharp\flat}:=prod(M_{\natural\sharp\flat})=prod(M_{\natural})prod(M_{\sharp})prod(M_{\flat})=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}.$
For later reference, we define supplementary elements $\lambda_{l}$ $(l\in\{1,\ldots,n\})$, $\lambda_{\sharp}(\mathfrak{j})$, $\gamma_{\sharp}(\mathfrak{j})$ $(\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\})$, $\lambda_{\flat}(a)$, $\gamma_{\flat}(a)$ $(a\in\{1,\ldots,r\})$, $\gamma_{\natural\sharp\flat}[d]\in A_{\Gamma}$ $(d\in\{1,\ldots,(1+r+\mathfrak{m})n\})$. For every $l\in\{1,\ldots,n\}$, we define $\lambda_{l}\in A_{V_{\ast}}\subset A_{\Gamma}$ by
(4.29) $\lambda_{l}:=prod(M_{\natural}[\ ,l])=v_{1,l}v_{2,l}\cdots v_{i,l}\cdots v_{k,l}.$
For every $\mathfrak{j}\in\{1,\ldots,\mathfrak{m}\}$, we define $\lambda_{\sharp}(\mathfrak{j}),\gamma_{\sharp}(\mathfrak{j})\in A_{\Gamma}$ by
(4.30) $\lambda_{\sharp}(\mathfrak{j})$ $:=prod(M_{\sharp}(\mathfrak{j})[\ ,l_{\sharp}(w_{\mathfrak{j}})]),$
(4.31) $\gamma_{\sharp}(\mathfrak{j})$ $:=prod(M_{\sharp}(\mathfrak{j})).$
For every $a\in\{1,\ldots,r\}$, we define $\lambda_{\flat}(a),\gamma_{\flat}(a)\in A_{\Gamma}$ by
(4.32) $\lambda_{\flat}(a)$ $:=prod(M_{\flat}(a)[\ ,l_{\flat}(i_{a},i_{a+1})]),$
(4.33) $\gamma_{\flat}(a)$ $:=prod(M_{\flat}(a)).$
Note that $\lambda_{\sharp}(\mathfrak{j})$ and $\lambda_{\flat}(a)$ can be regarded as modifications of $\lambda_{l_{\sharp}(w_{\mathfrak{j}})}$ and $\lambda_{l_{\flat}(i_{a},i_{a+1})}$, respectively. Also note that $\gamma_{\sharp}(\mathfrak{j})$ and $\gamma_{\flat}(a)$ can be regarded as modifications of $\gamma_{\natural}$. For every $d\in\{1,\ldots,(1+\mathfrak{m}+r)n\}$, let $\gamma_{\natural\sharp\flat}[d]$ be the product of the components of the first $d$ columns of $M_{\natural\sharp\flat}\in\mathcal{M}(k,(1+\mathfrak{m}+r)n;A_{\Gamma})$. That is,
(4.34) $\gamma_{\natural\sharp\flat}[d]:=prod(M_{\natural\sharp\flat}[\ ,1:d]).$
We recall the following for convenience:
  1. (1)
    $k$ is the number of indecomposable factors of the unique join decomposition of $\Gamma_{\ast}$;
  2. (2)
    $r$ is the length of the closed path (4.6) on $T$;
@@
Then, $U_{l}$ spans a clique containing $\Gamma_{0}$ in $\Gamma$; that is, $U_{l}\in\Delta_{\Gamma,\Gamma_{0}}$. Hence, we have a $k$-dimensional cube $[A_{V_{0}},A_{U_{l}}]$ in $C_{\Gamma,\Gamma_{0}}$. The hyperplanes dual to edges of $[A_{V_{0}},A_{U_{l}}]$ are of $v_{i,l}$-type ($i\in\{1,\ldots,k\}$). We denote such hyperplanes as $H_{i,l}$ ($i\in\{1,\ldots,k\}$).